OpenStudy (anonymous):

the sum of two numbers is 45 and the difference is 17 what are the numbers

OpenStudy (anonymous):

let two no be x and y x+y=45 x-y=17 solve thse to get no .

OpenStudy (anonymous):

let xand y be the numbers since they are unknown their sum which is x+y=45-----equation 1 their difference is x-y=17------equation2 from equation 2 x-y=17 x=17+y substitute xinyto equation 1 17+y+y=45 17+2y=45 make y the subject 2y=45-17 y=28/2 y=14 since x=17+y x=17+14 x=31

OpenStudy (texaschic101):

x + y = 45 ---> x = 45 - y x - y = 17 now sub 45 - y in for x in the second equation... x - y = 17 45 - y - y = 17 (combine like terms) 45 - 2y = 17 45 - 17 = 2y 28 = 2y 14 = y now sub 14 in for y in either of the equations... x + y = 45 x + 14 = 45 x = 45 - 14 x = 31 check... x + y = 45 31 + 14 = 45 45 = 45 (correct) check by the other equation... x - y = 17 31 - 14 = 17 17 = 17 (correct) ANSWER : The two numbers are 31 and 14